An Am broadcast station in Oman is allowed to transmit modulating frequencies up to 5 KHz. Calculate the maximum and minimum upper and lower sidebands and the total bandwidth occupied by the AM station if its transmitting frequency is 740 KHz.


1
Expert's answer
2021-12-08T10:00:41-0500

(a) We can find the maximum upper sideband as follows:


"f_{USB}=f_c+f_m=740\ KHz+5\ KHz=745\ KHz."

(b) We can find the minimum lower sideband as follows:


"f_{LSB}=f_c-f_m=740\ KHz-5\ KHz=735\ KHz."

(c) We can find the total bandwidth occupied by the AM station as follows:


"BW=2\times f_m=2\times5\ KHz=10\ KHz."